Find all points (x, y) where the functions f(x), g(x), h(x) have the same value:
f(x) = 2^x−5 + 3, g(x) = 2x − 5, h(x) = 8/x+ 10
1
Expert's answer
2020-11-16T08:18:43-0500

"f(x)=2^{x-5}+3,\ \ g(x)=2x-5, \ \ h(x)=8/x+10" .

Let us find all points "(x,y)" where functions "g(x)" and "h(x)" have the same value.

So, "g(x)=h(x)=y" .

"2x-5=8\/x+10" (multiply both sides of equation by "x" )

"2x^2-5x=8+10x"

"2x^2-15x-8=0"

"x_{1,2}=\frac{15\pm \sqrt{15^2-4\times 2\times (-8)}}{2\times 2}=\frac{15\pm17}{4}"

"x_1=8,\ \ x_2=-1/2"

Now we calculate "y_1=g(x_1)=11" , "y_2=g(x_2)=-6" .

We have two points "(8,\ 11)" and "(-1/2,\ -6)" where functions "g(x)" and "h(x)" have the same value.

Let is find which of these points satisfy the condition, that functions "f(x)" , "g(x)" and "h(x)" have the same value.

"f(x_1)=2^3+3=11=y_1"

"f(x_2)=2^{-1/2-5}+3=\frac{1}{32\sqrt{2}}+3\cancel{=}-6=y_2" .

Answer: "(8,\ 11)" .